# On the Identifying Power of Generalized Monotonicity

### v2

@@ A.3.2 Proof of the Theorem
$\begin{split}\sum_{r_{-j[k]}^{o}}q^{*}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})&=\sum_{r_{-j[k]}^{o}}q^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})\\ &=(1-\alpha_{j[k]})\sum_{r^{o}_{j[k]}\in\mathcal{Y}}\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})\\ &=(1-\alpha_{j[k]})\sum_{r^{o}}q(r^{o},r^{t})~{},\\ \sum_{r_{-j[k]}^{o}}q^{*}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})&=0\qquad\mbox{for}~{}y\not\in\{y^{L},y^{U}\}~{},\\ \sum_{r_{-j[k]}^{o}}q^{*}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})&=\sum_{r_{-j[k]}^{o}}q^{*}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})\\ &=\alpha_{j[k]}\sum_{r^{o}_{j[k]}\in\mathcal{Y}}\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})\\ &=\alpha_{j[k]}\sum_{r^{o}}q(r^{o},r^{t})~{}.\end{split}$ (24)
Set
$q^{*}(r^{o},r^{t})=q^{*}_{K(r^{t})}(r^{o},r^{t})\qquad\forall~{}r^{o}~{}.$
With this construction, $q^{*}(r^{o},r^{t})$ is non-negative and
$\sum_{r^{o}}q^{*}(r^{o},r^{t})=\sum_{r^{o}}q(r^{o},r^{t})~{}~{}\mbox{for all}~{}~{}r^{t}~{}.$
Note that, since Assumption 2.2 holds for $Q$, $r^{o}_{d}$ is a component of $r_{c}^{o}(r^{t})$ for $r^{t}$ such that $r^{t}_{z^{*}(d)}=d$, while $r^{o}_{d}$ is a component of $r_{n}^{o}(r^{t})$ for $r^{t}$ such that $r^{t}_{z^{*}(d)}\neq d$. Therefore, with this construction, for each $r^{t}$, $q^{*}(r^{o}_{c}(r^{t}),r^{t})=q(r^{o}_{c}(r^{t}),r^{t})~{}~{}\forall~{}r^{o}_{c}(r^{t})$ because of (24), and hence, by Lemma A.1, $q^{*}$ satisfies (20) so that $Q^{\ast}\in\mathbf{Q}_{0}(P,\mathbf{Q})$. That $q^{*}$ satisfies (20) implies that, for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{*}}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}=d\}]=\beta_{d|z^{*}(d)}~{}.$
We also have that for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{\ast}}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]$ $=\sum_{y\in\mathcal{Y}}~{}~{}\sum_{r^{t}:r^{t}_{z^{*}(d)}\neq d}~{}~{}\sum_{r^{o}:r^{o}_{d}=y}~{}~{}y~{}q^{*}(r^{o},r^{t})$
$=\sum_{y\in\mathcal{Y}}~{}~{}\sum_{r^{t}:r^{t}_{z^{*}(d)}\neq d}~{}~{}\sum_{r^{o}_{-d}}~{}~{}y~{}q^{*}((r^{o}_{-d},r^{o}_{d}=y),r^{t})$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})\sum_{r^{t}:r^{t}_{z^{*}(d)}\neq d}~{}~{}\sum_{r^{o}}q(r^{o},r^{t})$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})~{}Q\{D_{z^{*}(d)}\neq d\}$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})(1-\sum_{y}p_{yd|z^{*}(d)})~{},$
@@
where the third equality is using that (24) holds for $r^{t}$ such that $r^{t}_{z^{*}(d)}\neq d$, so that $d=j[k^{\prime}]$ for some $k^{\prime}$ in constructing $q^{*}(\cdot,r^{t})$ for that $r^{t}$, and the last equality is using that $Q$ satisfies (20). Thus, for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{*}}[Y_{d}]$ $=\mathbb{E}_{Q^{*}}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}=d\}]+\mathbb{E}_{Q^{*}}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}\neq d\}]$
$=\beta_{d|z^{*}(d)}+(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})(1-\sum_{y}p_{yd|z^{*}(d)})~{}.$
@@
For any $\theta$ contained in (5), we can thus chose $\alpha=(\alpha_{0},\alpha_{1},...,\alpha_{|\mathcal{D}|-1})^{\prime}\in[0,1]^{|\mathcal{D}|}$ such that $\theta({Q^{\ast}})=\theta$.
@@
$\mathbb{E}_{Q}[Y_{d}]=\mathbb{E}_{Q}[Y_{d}|Z=z]$ $=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D=d\}|Z=z]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}|Z=z]$
$=\mathbb{E}_{Q}[Y\mathbbm{1}\{D=d\}|Z=z]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}|Z=z]$
$=\beta_{d|z}+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}|Z=z]$
$\leq\beta_{d|z}+y^{U}P\{D\neq d|Z=z\}$
@@
Suppose $Q\in\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{*})\neq\emptyset$. Consider the upper endpoints of (5) and (7). For any $d\in\mathcal{D}$, $z\in\mathcal{Z}$,
$\mathbb{E}_{P}[Y$ $\mathbbm{1}\{D=d\}]\mid Z=z^{*}(d)]+y^{U}\mathbb{E}_{P}[1-\mathbbm{1}\{D=d\}\mid Z=z^{*}(d)]$
@@
$=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{*}(d)}=d\}]+y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z^{*}(d)}=d\}]-\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z}=d\}]-y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z}=d\}]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})(\mathbbm{1}\{D_{z^{*}(d)}=d\}-\mathbbm{1}\{D_{z}=d\})]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})(\mathbbm{1}\{D_{z^{*}(d)}=d,D_{z}\neq d\}-\mathbbm{1}\{D_{z^{*}(d)}\neq d,D_{z}=d\})]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})\mathbbm{1}\{D_{z^{*}(d)}=d,D_{z}\neq d\}]$
@@
where the first equality uses Assumption 2.1 and the fourth equality uses that $Q\{D_{z^{*}(d)}\neq d,D_{z}=d\}=0$ for all $Q$ satisfying Assumption 2.2. Since this inequality holds for all $z\in\mathcal{Z}$, we have that the upper endpoint of the interval in (5) is weakly smaller than the upper endpoint of the interval in (7). Conversely, the upper endpoint of (5) is contained in the set over which the upper endpoint of (7) is minimizing over, and thus the upper endpoint of (5) is weakly larger than the upper endpoint of (7). We conclude that the upper endpoints are the same. Parallel arguments show equivalence of the lower endpoints.
@@
Recall that Assumption 2.1 holds for $Q\in\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{*})$. Note that Assumption 2.1 implies Assumption 3.1, and hence $\mathbf{Q}_{1,2}^{\ast}\subseteq\mathbf{Q}_{1}^{\ast}\subseteq\mathbf{Q}_{3}^{\ast}$. As a result,
$\Theta_{0}(P,\mathbf{Q}_{1,2}^{*})\subseteq\Theta_{0}(P,\mathbf{Q}_{1}^{*})\subseteq\Theta_{0}(P,\mathbf{Q}_{3}^{*})~{}.$
By Lemma 3.2, when $\mathbf{Q}_{0}(P,\mathbf{Q}_{1,2}^{*})\neq\emptyset$, we have
$\Theta_{0}(P,\mathbf{Q}_{1,2}^{*})=\Theta_{0}(P,\mathbf{Q}_{3}^{*})~{}.$
@@ Appendix B Details of Examples
$Q\{D_{z^{*}(0)}\neq 0,~{}D_{z^{\prime}}=0~{}\mbox{for some}~{}z^{\prime}\neq z^{*}(0)\}=~{}Q\Big{\{}\bigcup_{d^{*}\neq 0,z^{\prime}\neq z^{*}(0)}D_{z^{*}(0)}=d^{*},~{}D_{z^{\prime}}=0\Big{\}}~{}.$
We wish to investigate whether there exists some $z^{*}(0)\in\mathcal{Z}$ such that the above probability is zero. Consider $z^{*}(0)$ equal to any value in $\mathcal{Z}\setminus\mathcal{Z}^{\dagger}$, i.e., any value of the instrument that does not target any treatment. For any fixed $d^{*}\neq 0,z^{\prime}\neq z^{*}(0)$, consider the event $\{D_{z^{*}(0)}=d^{*},~{}D_{z^{\prime}}=0\}$. Since $z^{*}(0)$ does not target any treatment and thus does not target $d^{*}$, $D_{z^{*}(0)}=d^{*}$ implies
$U_{0}-U_{d^{*}}\leq\underline{U}(d^{*})~{}.$ (27)
If $z^{\prime}$ targets $d^{*}$, then $D_{z^{\prime}}=0$ implies
$U_{0}-U_{d^{*}}\geq\overline{U}(d^{*})~{}.$ (28)
If $z^{\prime}$ does not target $d^{*}$, then $D_{z^{\prime}}=0$ implies
$U_{0}-U_{d^{*}}\geq\underline{U}(d^{*})~{}.$ (29)
Thus, the event $\{D_{z^{*}(0)}=d^{*},~{}D_{z^{\prime}}=0\}$ either requires (27) and (28) to jointly hold, which is a contradiction since $\overline{U}(d^{*})>\underline{U}(d^{*})$ , or requires (27) and (29) to jointly hold, which is a zero probability event given our assumption that the distribution of $(U_{d}:d\in\mathcal{D})$ is absolutely continuous w.r.t. Lebesgue measure. Thus $Q\{D_{z^{*}(0)}\neq 0,~{}D_{z^{\prime}}=0~{}\mbox{for some}~{}z^{\prime}\neq z^{*}(0)\}$ is a probability of a finite union of zero probability events, and thus, by Boole’s inequality, equals zero so that (3) holds for $d=0$. A parallel argument shows that (3) holds for any non-targeted treatment, and related argument shows that (3) holds for any targeted treatment. Thus, under the strict one-to-one targeting of Lee and Salanié (2023), when there are more values of the instrument than targeted treatments, Assumption 2.2 holds even though (11) is violated for some $d\in\mathcal{D}$.
@@
In contrast, the identified set that follows from imposing Assumption 2.1 alone, $\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$, is shown by Balke and Pearl (1997) to be
@@
It follows from $\mathbf{Q}\subseteq\mathbf{Q}_{1}^{\ast}$ that $\Theta_{0}(P,\mathbf{Q})\subseteq\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$, and thus (30) is contained in (31)–(32).
Next, we show that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, $\Theta_{0}(P,\mathbf{Q})$ is not given by (5) and $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$. We do so by providing a numerical example. Consider the $P$ specified in Table 1 and the $Q$ specified in Table 2, where we write $q(y_{0}y_{1},d_{0}d_{1})=Q\{Y_{d}=y_{d},D_{z}=d_{z},~{}(d,z)\in\mathcal{D}\times\mathcal{Z}\}$ and omit any $q(\cdot)=0$. One can check that $Q\in\mathbf{Q}$ and rationalizes $P$, so that $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. In this example, $Q\{Y_{0}=1\}=0.4274$, and thus the identified set for $Q\{Y_{0}=1\}$ relative to $\mathbf{Q}$ is the singleton $\{0.4274\}$. In contrast, evaluating (31) at $P$ gives the identified set for $Q\{Y_{0}=1\}$ relative to $\mathbf{Q}_{1}^{*}$ as $[0.3336,0.5212]$. We thus conclude that $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{1}^{\ast})$ for some $P$ that can be rationalized by $Q\in\mathbf{Q}$. Now consider evaluating the bounds of (5) at the same $P$. Doing so results in bounds on $Q\{Y_{0}=1\}$ given by $[0.1618,0.5445]$ if setting $z^{*}(0)=0$ and given by $[0.2656,0.8829]$ if setting $z^{*}(0)=1$. Therefore, no matter $z^{*}(0)=0$ or $1$, the bounds (5) is not the identified set for $Q\{Y_{0}=1\}$ relative to either $\mathbf{Q}_{1}^{*}$ or $\mathbf{Q}$.
@@
We will show that for some $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, we have $\Theta_{0}(P,\mathbf{Q})$ strictly smaller than (5) and $\Theta_{0}(P,\mathbf{Q}^{*}_{1})$. For this purpose we are only concerned with the validity of (33) instead of its sharpness. For the lower bounds, first note for $z\in\mathcal{Z}$,
$Q\{Y_{1}=1\}=Q\{Y_{1}=1|Z=z\}\geq Q\{Y_{1}=1,D=1|Z=z\}=Q\{Y=1,D=1|Z=z\}~{},$

### v3

@@ A.3.2 Proof of the Theorem
$\sum_{r_{-j[k]}^{o}}q^{\ast}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})$ $=\sum_{r_{-j[k]}^{o}}q^{\ast}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{L}),r^{t})$
$=(1-\alpha_{j[k]})\sum_{r^{o}_{j[k]}\in\mathcal{Y}}\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})$
$=(1-\alpha_{j[k]})\sum_{r^{o}}q(r^{o},r^{t})~{},$
$\sum_{r_{-j[k]}^{o}}q^{\ast}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y),r^{t})$ $=0\qquad\mbox{for}~{}y\not\in\{y^{L},y^{U}\}~{},$
$\sum_{r_{-j[k]}^{o}}q^{\ast}_{K(r^{t})}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})$ $=\sum_{r_{-j[k]}^{o}}q^{\ast}_{k}((r^{o}_{-j[k]},r^{o}_{j[k]}=y^{U}),r^{t})$
$=\alpha_{j[k]}\sum_{r^{o}_{j[k]}\in\mathcal{Y}}\sum_{r_{-j[k]}^{o}}q((r^{o}_{-j[k]},r^{o}_{j[k]}),r^{t})$
$=\alpha_{j[k]}\sum_{r^{o}}q(r^{o},r^{t})~{}.$ (18)
These equations state that under $q^{\ast}_{K(r^{t})}$, the constructed distribution in the final step $K(r^{t})$, the probability that $r^{o}_{j[k]}=y$ for each $r^{t}$ is zero if $y$ is not $y^{L}$ or $y^{U}$, is $\alpha_{j[k]}$ times the true probability of $r^{t}$ under $q$ if $y=y^{U}$, and is $1-\alpha_{j[k]}$ times the true probability of $r^{t}$ under $q$ if $y=y^{L}$.
Finally, set
$q^{\ast}(r^{o},r^{t})=q^{\ast}_{K(r^{t})}(r^{o},r^{t})\qquad\forall~{}r^{o}~{}.$
Part 2: verify property (a), $Q^{\ast}\in\mathbf{Q}$
With this construction, $q^{\ast}(r^{o},r^{t})$ is non-negative and from (18) we have
$\sum_{r^{o}}q^{\ast}(r^{o},r^{t})=\sum_{r^{o}}q(r^{o},r^{t})~{}~{}\mbox{for all}~{}~{}r^{t}~{}.$
Because we assume $\mathbf{Q}|_{Y}$ satisfies Assumption 3.1 and $Q\in\mathbf{Q}$, this implies that $Q^{\ast}\in\mathbf{Q}$.
Part 3: verify property (b), $P=Q^{\ast}T^{-1}$
Furthermore, for each $r^{t}$ and for all $r^{o}_{c}(r^{t})$, the construction of $q^{\ast}$ in (16) implies
$q^{\ast}(r^{o}_{c}(r^{t}),r^{t})$ $=\sum_{r^{o}_{n}(r^{t})\in\mathcal{Y}^{|\mathcal{N}(r^{t})|}}q^{\ast}(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$
$=\sum_{r^{o}_{j[1]}\in\mathcal{Y}}\cdots\sum_{r^{o}_{j[K(r^{t})]}\in\mathcal{Y}}q_{K(r^{t})}^{\ast}((r^{o}_{j[1]},\dots,r^{o}_{j[K(r^{t})]}),r^{o}_{c}(r^{t}),r^{t})$
$=\sum_{r^{o}_{j[1]}\in\mathcal{Y}}\cdots\sum_{r^{o}_{j[K(r^{t})]}\in\{y^{L},y^{U}\}}q_{K(r^{t})}^{\ast}(r^{o}_{-j[K(r^{t})]},r^{o}_{j[K(r^{t})]},r^{t})$
$=\sum_{r^{o}_{j[1]}\in\mathcal{Y}}\cdots\sum_{r^{o}_{j[K(r^{t})]}\in\mathcal{Y}}q_{K(r^{t})-1}^{\ast}(r^{o}_{-j[K(r^{t})]},r^{o}_{j[K(r^{t})]},r^{t})$
$=\sum_{r^{o}_{j[1]}\in\mathcal{Y}}\cdots\sum_{r^{o}_{j[K(r^{t})]}\in\mathcal{Y}}q_{0}^{\ast}((r^{o}_{j[1]},\dots,r^{o}_{j[K(r^{t})]}),r^{o}_{c}(r^{t}),r^{t})$
$=\sum_{r^{o}_{n}(r^{t})\in\mathcal{Y}^{|\mathcal{N}(r^{t})|}}q(r^{o}_{n}(r^{t}),r^{o}_{c}(r^{t}),r^{t})$
$=q(r^{o}_{c}(r^{t}),r^{t})~{}.$
Therefore for each $r^{t}$, $q^{\ast}(r^{o}_{c}(r^{t}),r^{t})=q(r^{o}_{c}(r^{t}),r^{t})~{}~{}\forall~{}r^{o}_{c}(r^{t})$, and hence by Lemma A.1, $q^{\ast}$ satisfies (14) so that $P=Q^{\ast}T^{-1}$. Thus $Q^{\ast}\in\mathbf{Q}_{0}(P,\mathbf{Q})$.
Part 4: verify property (c), $\theta(Q^{\ast})=\theta_{0}$
Note for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{\ast}}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]=\mathbb{E}_{P}[Y\mathbbm{1}\{D=d\}\mid Z=z^{\ast}(d)]=\beta_{d|z^{\ast}(d)}~{}.$
Further note since Assumption 2.2 holds for $Q$, we have that for each $d\in\mathcal{D}$ and for each $r^{t}$ if $r^{t}_{z^{\ast}(d)}=d$ then $r^{o}_{d}$ is a component of $r_{c}^{o}(r^{t})$ and $d\in\mathcal{N}(r^{t})^{c}$, while if $r^{t}_{z^{\ast}(d)}\neq d$ then $r^{o}_{d}$ is a component of $r_{n}^{o}(r^{t})$ and $d\in\mathcal{N}(r^{t})$. Then we also have that for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{\ast}}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}\neq d\}]$ $=\sum_{y\in\mathcal{Y}}~{}~{}\sum_{r^{t}:r^{t}_{z^{\ast}(d)}\neq d}~{}~{}\sum_{r^{o}:r^{o}_{d}=y}~{}~{}y~{}q^{\ast}(r^{o},r^{t})$
$=\sum_{y\in\mathcal{Y}}~{}~{}\sum_{r^{t}:r^{t}_{z^{\ast}(d)}\neq d}~{}~{}\sum_{r^{o}_{-d}}~{}~{}y~{}q^{\ast}((r^{o}_{-d},r^{o}_{d}=y),r^{t})$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})\sum_{r^{t}:r^{t}_{z^{\ast}(d)}\neq d}~{}~{}\sum_{r^{o}}q(r^{o},r^{t})$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})~{}Q\{D_{z^{\ast}(d)}\neq d\}$
$=(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})(1-\sum_{y}p_{yd|z^{\ast}(d)})~{},$
@@
where the third equality is using that (18) holds for $r^{t}$ such that $r^{t}_{z^{\ast}(d)}\neq d$, so that $d=j[k^{\prime}]$ for some $k^{\prime}$ in constructing $q^{\ast}(\cdot,r^{t})$ for that $r^{t}$, and the last equality is using that $Q$ satisfies (14). Thus, for each $d\in\mathcal{D}$,
$\mathbb{E}_{Q^{\ast}}[Y_{d}]$ $=\mathbb{E}_{Q^{\ast}}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]+\mathbb{E}_{Q^{\ast}}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}\neq d\}]$
$=\beta_{d|z^{\ast}(d)}+(\alpha_{d}y^{U}+(1-\alpha_{d})y^{L})(1-\sum_{y}p_{yd|z^{\ast}(d)})~{}.$
@@
For any $\theta_{0}$ contained in (5), we can thus choose $\alpha=(\alpha_{0},\alpha_{1},...,\alpha_{|\mathcal{D}|-1})^{\prime}\in[0,1]^{|\mathcal{D}|}$ such that $\theta({Q^{\ast}})=\theta_{0}$.
@@
$\mathbb{E}_{Q}[Y_{d}]=\mathbb{E}_{Q}[Y_{d}\mid Z=z]$ $=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D=d\}\mid Z=z]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}\mid Z=z]$
$=\mathbb{E}_{Q}[Y\mathbbm{1}\{D=d\}\mid Z=z]+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}\mid Z=z]$
$=\beta_{d|z}+\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D\neq d\}\mid Z=z]$
$\leq\beta_{d|z}+y^{U}P\{D\neq d\mid Z=z\}$
@@
Suppose $Q\in\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$. Consider the upper endpoints of (5) and (8). For any $d\in\mathcal{D}$, $z\in\mathcal{Z}$,
$\mathbb{E}_{P}[Y$ $\mathbbm{1}\{D=d\}]\mid Z=z^{\ast}(d)]+y^{U}\mathbb{E}_{P}[1-\mathbbm{1}\{D=d\}\mid Z=z^{\ast}(d)]$
@@
$=\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]+y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z^{\ast}(d)}=d\}]-\mathbb{E}_{Q}[Y_{d}\mathbbm{1}\{D_{z}=d\}]-y^{U}\mathbb{E}_{Q}[1-\mathbbm{1}\{D_{z}=d\}]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})(\mathbbm{1}\{D_{z^{\ast}(d)}=d\}-\mathbbm{1}\{D_{z}=d\})]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})(\mathbbm{1}\{D_{z^{\ast}(d)}=d,D_{z}\neq d\}-\mathbbm{1}\{D_{z^{\ast}(d)}\neq d,D_{z}=d\})]$
$=\mathbb{E}_{Q}[(Y_{d}-y^{U})\mathbbm{1}\{D_{z^{\ast}(d)}=d,D_{z}\neq d\}]$
@@
where the first equality uses Assumption 2.1 and the fourth equality uses that $Q\{D_{z^{\ast}(d)}\neq d,D_{z}=d\}=0$ for all $Q$ satisfying Assumption 2.2. Since this inequality holds for all $z\in\mathcal{Z}$, we have that the upper endpoint of the interval in (5) is weakly smaller than the upper endpoint of the interval in (8). Conversely, the upper endpoint of (5) is contained in the set over which the upper endpoint of (8) is minimizing over, and thus the upper endpoint of (5) is weakly larger than the upper endpoint of (8). We conclude that the upper endpoints are the same. Parallel arguments show the equivalence of the lower endpoints.
@@
[…]
By Lemma 3.2, when $\mathbf{Q}_{0}(P,\mathbf{Q}_{E,M}^{\ast})\neq\emptyset$, we have
[…]
@@ Appendix B Details of Examples
$Q\{D_{z^{\ast}(0)}\neq 0,~{}D_{z^{\prime}}=0~{}\mbox{for some}~{}z^{\prime}\neq z^{\ast}(0)\}=~{}Q\left\{\bigcup_{d^{\ast}\neq 0,z^{\prime}\neq z^{\ast}(0)}D_{z^{\ast}(0)}=d^{\ast},~{}D_{z^{\prime}}=0\right\}~{}.$
We wish to investigate whether there exists some $z^{\ast}(0)\in\mathcal{Z}$ such that the above probability is zero. Consider $z^{\ast}(0)$ equal to any value in $\mathcal{Z}\setminus\mathcal{Z}^{\dagger}$, i.e., any value of the instrument that does not target any treatment. For any fixed $d^{\ast}\neq 0,z^{\prime}\neq z^{\ast}(0)$, consider the event $\{D_{z^{\ast}(0)}=d^{\ast},~{}D_{z^{\prime}}=0\}$. Since $z^{\ast}(0)$ does not target any treatment and thus does not target $d^{\ast}$, $D_{z^{\ast}(0)}=d^{\ast}$ implies
$U_{0}-U_{d^{\ast}}\leq\underline{U}(d^{\ast})~{}.$ (21)
If $z^{\prime}$ targets $d^{\ast}$, then $D_{z^{\prime}}=0$ implies
$U_{0}-U_{d^{\ast}}\geq\overline{U}(d^{\ast})~{}.$ (22)
If $z^{\prime}$ does not target $d^{\ast}$, then $D_{z^{\prime}}=0$ implies
$U_{0}-U_{d^{\ast}}\geq\underline{U}(d^{\ast})~{}.$ (23)
Thus, the event $\{D_{z^{\ast}(0)}=d^{\ast},~{}D_{z^{\prime}}=0\}$ either requires (21) and (22) to jointly hold, which is a contradiction since $\overline{U}(d^{\ast})>\underline{U}(d^{\ast})$ , or requires (21) and (23) to jointly hold, which is a zero probability event given our assumption that the distribution of $(U_{d}:d\in\mathcal{D})$ is absolutely continuous w.r.t. Lebesgue measure. Thus $Q\{D_{z^{\ast}(0)}\neq 0,~{}D_{z^{\prime}}=0~{}\mbox{for some}~{}z^{\prime}\neq z^{\ast}(0)\}$ is a probability of a finite union of zero probability events, and thus, by Boole’s inequality, equals zero so that (3) holds for $d=0$. A parallel argument shows that (3) holds for any non-targeted treatment, and related argument shows that (3) holds for any targeted treatment. Thus, under the strict one-to-one targeting of Lee and Salanié (2023), when there are more values of the instrument than targeted treatments, Assumption 2.2 holds even though (11) is violated for some $d\in\mathcal{D}$.
@@
In contrast, the identified set that follows from imposing Assumption 2.1 alone, $\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$, is shown by Balke and Pearl (1997) to be
@@
It follows from $\mathbf{Q}\subseteq\mathbf{Q}_{E}^{\ast}$ that $\Theta_{0}(P,\mathbf{Q})\subseteq\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$, and thus (24) is contained in (25)–(26).
Next, we show that there exists a $P$ for which $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, $\Theta_{0}(P,\mathbf{Q})$ is not given by (5) and $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$. We do so by providing a numerical example. Consider the $P$ specified in Table 1 and the $Q$ specified in Table 2, where we write $q(y_{0}y_{1},d_{0}d_{1})=Q\{Y_{d}=y_{d},D_{z}=d_{z},~{}(d,z)\in\mathcal{D}\times\mathcal{Z}\}$ and omit any $q(\cdot)=0$. One can check that $Q\in\mathbf{Q}$ and rationalizes $P$, so that $Q\in\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$. In this example, $Q\{Y_{0}=1\}=0.4274$, and thus the identified set for $Q\{Y_{0}=1\}$ relative to $\mathbf{Q}$ is the singleton $\{0.4274\}$. In contrast, evaluating (25) at $P$ gives the identified set for $Q\{Y_{0}=1\}$ relative to $\mathbf{Q}_{E}^{\ast}$ as $[0.3336,0.5212]$. We thus conclude that $\Theta_{0}(P,\mathbf{Q})\subsetneqq\Theta_{0}(P,\mathbf{Q}_{E}^{\ast})$ for some $P$ that can be rationalized by $Q\in\mathbf{Q}$. Now consider evaluating the bounds of (5) at the same $P$. Doing so results in bounds on $Q\{Y_{0}=1\}$ given by $[0.1618,0.5445]$ if setting $z^{\ast}(0)=0$ and given by $[0.2656,0.8829]$ if setting $z^{\ast}(0)=1$. Therefore, no matter $z^{\ast}(0)=0$ or $1$, the bounds (5) is not the identified set for $Q\{Y_{0}=1\}$ relative to either $\mathbf{Q}_{E}^{\ast}$ or $\mathbf{Q}$.
@@
We will show that for some $P$ such that $\mathbf{Q}_{0}(P,\mathbf{Q})\neq\emptyset$, we have $\Theta_{0}(P,\mathbf{Q})$ strictly smaller than (5) and $\Theta_{0}(P,\mathbf{Q}^{\ast}_{1})$. For this purpose we are only concerned with the validity of (27) instead of its sharpness. For the lower bounds, first note for $z\in\mathcal{Z}$,
$Q\{Y_{1}=1\}=Q\{Y_{1}=1\mid Z=z\}\geq Q\{Y_{1}=1,D=1\mid Z=z\}=Q\{Y=1,D=1\mid Z=z\}~{},$